How many atoms are in 0.340 moles of sodium?


1
Expert's answer
2021-02-15T03:09:11-0500

1 mole = 6.02 * 10^23 particles

0.340 moles Na * (6.02 * 10^23 atoms) / 1mol

= 2.05 * 10^23 atoms Na

= 2.05 * 10^23 atoms Sodium